Grade 10Analytical Geometry

The length of the side of an equilateral triangle inscribed in a circle of radius 8cm is....

Anand kumar
9 Years agoGrade 10

1 Answer

Saurabh Koranglekar
6 Years ago
Dear student

The radius is 2/3 of the height

ie h =(3/2)*r
= 12

using trigonometry
h/(side of eq triangle ) = sin (60)

so the side can be found after simplifying this

Regards